Parsons Company is planning to produce 2,000 units of product in 2020. Each unit requires 1.40 pounds of materials at $7.60 per pound and a half-hour of labor at $16.00 per hour. The overhead rate is 80% of direct labor.

(a) Compute the budgeted amounts for 2020 for direct materials to be used, direct labor, and applied overhead.

Direct materials $ enter a dollar amount

Direct labor $ enter a dollar amount

Overhead $ enter a dollar amount

(b) Compute the standard cost of one unit of product. (Round answer to 2 decimal places, e.g. 2.75.)

Standard cost $ enter the standard cost of one unit of product in dollars rounded to 2 decimal places
Answer #1

a) Budgeted amount

Direct material = 2000*1.4*7.6 = 21280

Direct labor = 2000*.50*16 = 16000

Overhead cost = 16000*80% = 12800

b) Standard cost per one unit = (21280+16000+12800)/2000 = 25.04

> How did you get .50 to calculate the direct labor?

Anthony Arjune Mon, Nov 15, 2021 9:02 AM